OpenStudy (anonymous):

Write the equation of the line {y=mx+b} 1)Given:m=0&b=8 2)Given: the slope is undefined And the line passes through (4,3)

OpenStudy (anonymous):

1) y=mx+b y=0x+8 y=8 ~ equation of the line. 2)y=mx+b 3=m(4) + 8 3-8=4m -5=4m -5/4=m ~ y=-5/4x+8 ~ equation of line